Score: /12 Attempts: 11. Problem 10-08 eBook Problem 10-08 NPVS, IRRs, and MIRRs for Independent Projects Edelman Engineering is considering including two pieces of equipment, a truck and an overhead pulley system, in tax cash flows, including depreciation, are as follows: Pulley $7,500 Year Truck $5,100 5,100 5,100 5,100 5,100 7,sv0 7,500 7,500 7,500 a. Calculate the IRR for each project Round your answers to two decimal places. Truck: what is the correct accept/reject decision for this project? Pulley: What is the correct acceptireject decision for this project? b. Caloulate the NPV for each project. Round your answers to the nearest dollar, if necessa Truck: ry. Enter each answer as what is the correct accept/reject decision for this project?
system, in this years capital budget. The projects are independent. The cash outlay for the truck is $18,000 and that for the pull Enter each answer as a whole number, For example, do not enter 1,000,000 as 3 million
Q Search this course independent. The cash outlay for the truck is $18,000 and that for the pulley system is $22,000. The firms cost of capital is 14%. After. ple, do not enter 1,000,000 as 1 million
Problem Set 2 Pulley $7,500 7,500 7,500 7,500 7,500 Truck Year $5,100 2 5,100 5,100 5,100 5 5,100 a. Calculate the IRR for each project. Round your answers to two decimal places. Truck: What is the correct accept/reject decision for this project? Pulley: What is the correct accept/reject decision for this project? b. Calculate the NPy for each project. Round your answers to the nearest dollar, If necessary. Enter each answer as a whole Truck: $ what is the correct accept/reject decision for this project? Pulley: What is the correct accept/reject decision for this preject? Caloulate the MIR for each preject. Round your answers to two decimal places. Truck What is the correct accestireject decision for this project? 다
ulley: What is the correct accept/reject decision for this project? Select- b. Calculate the NPV for each project. Round your answers to the nearest dollar, if necessary. Enter ea ruck:$ What is the correct accept/reject decision for this project? Select Pulley:$ What is the correct accept/rej-ikt decision for this project? -Select- c. Calculate the MIRR for each project. Round your answers to two decimal places. Truck: What is the correct accept/reject decision for this project? Pulley What is the correct accept/reject decision for this project?
Answer #1

Working notes:

Calculation of IRR and NPV

Cumulative discount factor for 5 years @14% cost of capital= 3.433 (apprx.)

NPV(PV of cash flows after tax- total outlay) of:

1) truck= (3.433*5100)- 18000=$(491.7)

2) pulley= (3.433* 7500)- 22000= $3747.5

Cumulative discount factor at 10% for 5 years= 3.791

Cumulative discount factor at 22% for 5 years= 2.864

NPV of truck at 10% cost of capital= (3.791*5100)- 18000= $1334.1

NPV of poultr at 22% cost of capital= (2.864*7500)- 22000= $(520)

Computation of future value of cash flows

Cumulative reinvestment factor @14% for 5 years= 6.61

FV of cash flows:

  1. For Truck= 6.61*5100= $33711
  2. For pulley= 6.61*7500=$49575

Answer to question a.

If the IRR is higher than the expected rate of return(here, 14%), the project should be accepted. Otherwise, it should be rejected.

IRR (truck)= 14%- [[(14-10)* 491.7]/[1334.1-(491.7)]]= 12.923%(apprx)

The irr is less than 14%, hence, the other should be rejected.

IRR (pulley)= 14%+ [[(22-14)*3747.5]/[3747.5-(520)]]= 21.025% (apprx.)

The irr is greater than 14% and hence the project should be accepted.

Answer to question b.

The project should be accepted only if the Npv is positive.

NPV of truck=$(491.7) reject

NPV of pulley= $3747.5 accept

Answer to question c

Mirr had the same criteria as irr. The project should be selected only if the mirr is greater than the expected rate of return (here, 14%)

MIRR= [nth root of (FV of cash flows at reinvestment rate/PV of total cash outlay) ]-1

MIRR (truck)= [5th root ( 33711/ 18000)]-1= 0.134 ie, 13.4%

Mirr is less than 14%. The project should be rejected.

MIRR (pulley)= [5th root(49575/22000)]-1=0.176 ie, 17.6%

MIRR is greater than 14% and hence the project should be accepted.

Note: please refer to the working notes above for the calculations.